OpenStudy (neonumbrella5115):

Is 0 to the power of two equal to zero? Why or why not?

Parth (parthkohli):

\[0^2 = 0 \times 0\]What do you think \(0 \times 0 \) is?

zepdrix (zepdrix):

No no you must be thinking backwards :)\[\Large\rm 0^1\ne 1\]\[\Large\rm 1^0=1\]\[\Large\rm 5^0=1\]\[\Large\rm a^0=1,\qquad a\ne0\]

OpenStudy (neonumbrella5115):

@zepdrix : Yes, I had it mixed up. Thanks!